Cody is twice as old as Evan and seven years ago the sum of their age was 7. Find the age of each boy now.
Answer by lwsshak3(11628)

Cody is twice as old as Evan and seven years ago the sum of their age was 7. Find the age of each boy now.
let x=Evan's age now
2x=Cody's age now
7 yrs ago:
x-7=Evan's age
2x-7=Cody's age
x-7+2x-7=7
3x=21
x=7
2x=14
Evan's age now=7
Cody's age now=14